What will it cost to carpet a room with indoor/outdoor carpet if the room is 10 feet wide and 12 feet long? The carpet cost 12.51 per square yard.
Found 2 solutions by Zenock, w3suhl33:
Answer by Zenock(15)

There are 3 feet in a yard.
So 10 feet is 10/3 yards and 12 feet is 12/3 = 4 yards.
To find how many square yards you need 10/3 x 4 = 20/3 square yards.
at 12.51 per square yard you get 20/3 * 12.51 = 83.40.
Of course in the real world this would not be accurate because they will make you pay for a full yard regardless and you end up scrap left over because if the room is 10 feet by 12 feet and the carpet comes in bolts 11 feet wide you would have to buy a 11 feet by 12 feet piece of carpet, and you would have to pay for labor, and glue, and padding, and etc... etc... etc...

Answer by w3suhl33(2)

1) Convert the problem to yards (3 ft = 1 yd)


10 ft = 10/3 yd
12 ft = 12/3 = 4 yd


2) Find area of the room (length x width)


10/3 x 4 = 40/3 or 13 1/3 sq yd


3) Find total (area x cost per sq yd)


40/3 x 12.51 = 166.8


4) Answer


It will cost $166.80 to carpet a room if the room is 10 ft wide and 12 ft long